# Large Language Models Reflect Human Citation Patterns with a Heightened Citation Bias

### v2

@@ Abstract
Citation practices are crucial in shaping the structure of scientific knowledge, yet they are often influenced by contemporary norms and biases. The emergence of Large Language Models (LLMs) like GPT-4 introduces a new dynamic to these practices. Interestingly, the characteristics and potential biases of references recommended by LLMs that entirely rely on their parametric knowledge, and not on search or retrieval-augmented generation, remain unexplored. Here, we analyze these characteristics in an experiment using a dataset of 166 papers from AAAI, NeurIPS, ICML, and ICLR, published after GPT-4’s knowledge cut-off date, encompassing 3,066 references in total. In our experiment, GPT-4 was tasked with suggesting scholarly references for the anonymized in-text citations within these papers. Our findings reveal a remarkable similarity between human and LLM citation patterns, but with a more pronounced high citation bias in GPT-4, which persists even after controlling for publication year, title length, number of authors, and venue. Additionally, we observe a large consistency between the characteristics of GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. By analyzing citation graphs, we show that the references recommended by GPT-4 are embedded in the relevant citation context, suggesting an even deeper conceptual internalization of the citation networks. While LLMs can aid in citation generation, they may also amplify existing biases and introduce new ones, potentially skewing scientific knowledge dissemination. Our results underscore the need for identifying the model’s biases and for developing balanced methods to interact with LLMs in general. [^0] Data and code are available at https://github.com/AndresAlgaba/LLM_citation_patterns.
1 Introduction
@@
Figure 1: Overview of our experiment evaluating the characteristics and biases of the references generated by GPT-4, when tasked to suggest references for anonymized in-text citations. We collect $166$ papers from the cs.LG category on arXiv which are published in the main tracks of AAAI, NeurIPS, ICML, and ICLR, and only appeared available online after GPT-4-0613’s knowledge cut-off date. We split the main content, which includes the author information, conference information, abstract, and introduction, from the ground truth references. GPT-4 is prompted to generate suggestions of scholarly references for the anonymized in-text citations in the main content. We verify the existence of the generated references via Semantic Scholar and compare the characteristics, such as title length, publication year, venue, and number of authors, of the existing and non-existent generated references with the ground truth. For the existing generated references, we also compare additional characteristics, such as the number of citations and references, and analyze the properties of their citation networks.
@@
Large Language Models (LLMs) have revolutionized natural language understanding and generation, driving scientific research forward by assisting in all steps of the scientific process, ranging from identifying research gaps to accelerating complex data analysis [3, 10, 24, 28, 40, 44]. One particularly interesting application is the generation of suggestions for appropriate scholarly references [27, 35]. Yet, without the aid of web browsing or retrieval-augmented generation, these models rely entirely on their parametric knowledge encapsulated during their (pre-)training [5, 6, 14, 38, 41]. Our research focuses on this intrinsic citation behavior of GPT-4, exploring how the model recommends references based on its training data, and highlighting the potential biases that arise from this internalized knowledge [1, 23].
@@
Biases in citation practices have long been a subject of scrutiny in the scientific community [8, 29]. Common biases include a preference for recent publications [4], shorter titles [20], and high-profile publication venues [18]. Moreover, a well-documented phenomenon is the “Matthew effect,” where highly cited papers tend to accumulate even more citations [36]. The number of authors on a paper can also influence its likelihood of being cited, with solo and small author groups often cited less frequently than larger teams [9]. By examining how these biases manifest in the GPT-4 generated references, we aim to uncover the underlying patterns and their potential to amplify existing biases and introduce new ones, potentially skewing scientific knowledge dissemination.
In our experiment, we let GPT-4 suggest scholarly references for anonymized in-text citations within a paper and compare the characteristics and citation networks of the GPT-4 generated references against the ground truth. We provide a comprehensive analysis of $166$ papers which are published in the main tracks of AAAI, NeurIPS, ICML, and ICLR and only available online after GPT-4-0613’s knowledge cut-off date. We find that GPT-4 exhibits strong preferences for highly cited papers, which persists even after controlling for multiple confounding factors such as publication year, title length, venue, and number of authors. Additionally, we observe a large consistency between GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. By analyzing citation graphs, we show that the references recommended by GPT-4 are embedded in the relevant citation context, suggesting an even deeper conceptual internalization of the citation networks. While LLMs can aid in citation generation, we find they may also amplify existing biases and introduce new ones, potentially skewing scientific discourse. Our results underscore the need for identifying the model’s biases and for developing balanced methods to interact with LLMs in general [25].
@@
Our setting differs from previous work which either let the LLM generate short papers or literature reviews, or is prompted for the most important papers on a certain topic [35]. We argue that these methods are more susceptible to the LLM’s knowledge cut-off [11, 13, 26, 32, 33]. Moreover, the evaluation of the suggested references mostly focuses on their existence, bibliometric accuracy, or qualitative judgement by domain experts [27]. Finally, another strand of the literature focuses on improving LLMs via search and retrieval-augmented generation [21] or to reduce their hallucination rate via self-consistency [2] to enhance their capabilities in systematic literature reviews [31].
2 Generating Citations with Large Language Models
@@
Our data consists of $166$ papers published at AAAI ($25$), NeurIPS ($72$), ICML ($38$), and ICLR ($31$) for a total of $3,066$ references. Our data collection process is depicted in Figure 1 (see Appendix A for more details) and begins by retrieving relevant papers from arXiv, focusing on those within the machine learning category (cs.LG) and posted between March 2022 and October 2023 (after GPT-4-0613’s knowledge cut-off date). We ensure these papers reference the target conferences and exclude those that do not meet our criteria using specific keywords (e.g., workshop). The papers are also verified on Semantic Scholar where we store additional metadata, such as all the reference titles with corresponding Semantic Scholar IDs to construct the citation networks (see Appendix Table LABEL:tab:paper_details for a full list of all the included papers).
@@
We split the main content, which includes the author information, conference information, abstract, and introduction, from the ground truth references. Next, we prompt GPT-4 with the main content accompanied by the default system message: “You are a helpful assistant” and the following prompt:
@@
We then post-process GPT-4’s response to extract the title, venue, publication year, author names, and number of authors for each generated reference (see Appendix A for more details). To assess the robustness of this approach, we repeat this “vanilla” approach five times for all $166$ papers.
A well-known issue in text generation by LLMs are hallucinations or confabulations, which refer to generated content that is nonsensical or untruthful in relation to certain sources, i.e., factual mistakes about historical events [43]. This is particularly problematic for the generation of scholarly references, as LLMs can fabricate references that do not exist or introduce subtle errors, making it impossible to retrieve the actual references [35]. There are two main approaches to verify the existence of LLM-generated references: one involves asking additional questions to the LLM to verify its self-consistency [2], and the second approach utilizes external databases to verify a reference’s existence [7]. In our experiment, we opt for the latter and determine via title and author names matching with Semantic Scholar entries whether the generated references exist (see Appendix A for more details). Finally, we also build on our “vanilla” approach, by introducing an “iterative” approach where we continue to prompt GPT-4 and indicate which generated references do not exist and ask to replace those with existing ones (see Appendix A for more details). The previously existing generated and the newly generated references are then merged.
In Table 1, we report the summary statistics for each of the five vanilla (and iterative: results between brackets) runs. On average, $65\%$ ($86\%$) of the generated references match with an entry in Semantic Scholar, while $13\%$ ($14\%$) and $17\%$ ($20\%$) of them appear in the introduction or paper itself, respectively. We further show that about $7\%$ ($7\%$) of the generated and ground truth references match pairwise. However, this number grows to $13\%$ ($14\%$) if we only consider the uniquely identifiable references (i.e., omitting references included in [4–8] as there is no one-to-one correspondence).
Table 1: Summary statistics (numbers in % with respect to total number of references).
@@
Pairwise match for 7.0 7.2 6.3 6.9 6.7
all references (7.1) (7.3) (6.6) (7.0) (7.1)
Pairwise match for uniquely 12.5 13.7 12.5 13.7 13.3
@@
3 Reflecting Human Citation Patterns
Figure 2 displays the characteristics of the ground truth and generated references, and separately the characteristics of the generated references which match with a Semantic Scholar entry, and those which do not exist according to this database. Overall, we observe a remarkable similarity between human and LLM citation patterns and a large consistency between GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. In Appendix Figure B1, we also show that the newly generated papers from the “iterative” approach show nearly identical distributions.
The distributions of the title lengths show that the existing generated reference titles tend to be the shortest, while the non-existent generated reference titles are more similar in length to the ground truth, which indicates a learned pattern. Overall, the first effect dominates, so the average is skewed towards shorter titles for generated references (Figure 2b). The temporal analysis reveals a similar pattern where the non-existent generated references follow a distribution that is more similar to the ground truth than the existent ones (Figure 2c). The existent generated references are on average less recent, which might be understood as follows: the sources from which the references can be learned (before the knowledge cut-off) must be more recent than the source itself, and this takes some time.
@@ 3 Reflecting Human Citation Patterns
Most prominently, we observe a significant citation bias in the existing generated references, which have a median citation count of $1,326$ higher than ground truth references (Figure 2f). In Appendix Figure B3, we compare the characteristics for the corresponding ground truth references of existing and non-existent references, and for the existing references which also appear in the paper itself. We observe that the ground truth papers which correspond to existing references that appear in the paper itself have by far the most citations, followed by the existing references, and the ground truth papers corresponding to non-existent references have the lowest numbers of citations. These findings further indicate the tendency for GPT-4 to more easily generate references to highly cited papers. Finally, the distribution of references indicates that ground truth references cite slightly more papers than existing generated references, with a median difference of $6$ (Figure 2g).
Figure 2: Properties of the ground truth and generated introduction references for the vanilla strategy. This figure displays the properties of the ground truth ($n=14,554$, in blue) and generated references ($n=14,554$, in green), further subdividing the generated references into existing ($n=9,376$, in orange) and non-existent ($n=5,178$, in red), from the original data sources of five runs for the vanilla strategy with GPT-4. a, The average percentage of existing generated references in total ($64.4\%$) and for each publication venue under the vanilla and iterative strategy, with dots representing the percentage for each of the five runs with GPT-4. b, The distribution of the number of characters in the title shows some differences between the ground truth (median $62$) and generated (median $58$) references with the non-existent being slightly longer and with a larger variance compared to the existing generations. c, The distribution over time reveals that ground truth references are relatively more recent than generated references, with most references post-$2010$. The temporal distribution of the non-existent generated references aligns more with the ground truth than the existing generated references. d, The distribution of the number of authors demonstrates a disparity between the ground truth and generated references, having median values of three and two, respectively. However, GPT-4 more often generates “et al.” which does not allow for an exact computation, especially for the non-existent references. e, The distribution of publication venues shows that for most venues the ground truth has the highest relative representation, followed closely by existing references. The non-existent references deviate more from the ground truth as the proportion of “Others” is substantially larger. f, The distributions of citations for ground truth and existing generated references reveal a substantial citation bias in the generated references with a difference in median citations of $1,326$. g, Finally, the distribution of references shows that ground truth references cite slightly more papers than the existing generated references with a median difference in median references of $6$.
@@
Figure 3: The citation bias in existing generated references is not due to the recency of ground truth references. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher citation count compared to their corresponding ground truth ($n=9,376$, in blue) across subperiods. a, The citation counts across time for the ground truth and existing generated references reveal that the most recent references have a relatively low number of citations. The difference in median citations between the existing generated references and their corresponding ground truth references is 1,257. Since the ground truth references are relatively more recent compared to the existing generated references, we examine whether the observed citation bias is related to the recency of ground truth references. b, The distributions of citations by subperiod reveal that the existing generated references consistently exhibit a higher citation count than their corresponding ground truth counterparts. c, The difference in median citations is most pronounced in the early and late subperiods, i.e., $\leq 1988$, 2010-2016, and 2017-2023.
Figure 4: The citation bias in existing generated references is consistent across title length, number of authors, and publication venue. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher citation count compared to their corresponding ground truth ($n=9,376$, in blue) across title length, number of authors, and publication venue. a, The citation counts across the number of characters in title reveals that the discrepancy in number of citations between the existing generated and ground truth references is consistent over various title lengths. b and c, The distributions of citation counts per number of authors and publication venues show that the existing generated references consistently exhibit a higher citation count than their corresponding ground truth counterparts.
Figure 5: The generated references display similar citation network properties as the ground truth references but with a heightened citation bias. This figure displays how the existing generated references ($n=2945$, first run of vanilla strategy) are embedded in the citation network of the focal papers ($m=166$ in total). a, We depict the connections between the focal paper, the ground truth references, and the existing generated references by showing the underlying citation graphs. An arrow from A to B indicates that A cites B. We identify the focal paper (in blue), generated references that appear in the introduction (in green) or in the paper (in yellow), generated references that are linked to ground truth or other generated references (in orange), generated references that are completely isolated (in purple), and ground truth references that are not cited by GPT-4 (in gray). b, The majority of generated references does not appear in the introduction or paper itself, but is somehow connected to the ground truth references as only a small fraction of generated references is completely isolated. c, The heightened citation bias is most emphasized for generated references that appear in the introduction or the paper, with isolated generated references having the lowest number of citations. d, The normalized average clustering coefficients of the ground truth (green and grey nodes) and the existing generated references (green, yellow, orange, and purple nodes) indicate that GPT-4’s internalization of citation patterns extends to citation network properties. The clustering coefficient for a node A is given by #triangles through A#possible triangles through $\frac{\#\text{triangles through A}}{\#\text{possible triangles through A}}$. The average is computed across the coefficients of all nodes in the respective graph (excl. nodes with coefficient zero) and indicates the tendency of the respective references to appear in clusters. e, The non-isolated generated references are tightly connected to the ground truth references, both on an individual level (Boolean edge density) as well as an aggregate level (edge expansion). The Boolean edge density is the fraction of non-isolated generations (orange nodes) that are connected to at least one ground truth reference (green and grey nodes) per focal paper. The edge expansion between those two sets is defined as the number of edges between the two sets divided by the smallest set size. f, The number of references is similar across all categories, except for the isolated generated references which have substantially less references.
4 Heightened Citation Bias
@@
In Figure 4, we find that the heightened citation bias in generated references remains also after controlling for other possible confounding factors, such as title length, number of authors, and publication venue. Figure 4a shows that the citation discrepancy is consistent across different title lengths, with generated references tending to be more highly cited regardless of title length variations. Figure 4(b,c) examine the distributions of citation counts by the number of authors and publication venues, respectively. The data indicates that, regardless of the number of authors or the publication venue, generated references have higher citation counts compared to the ground truth references. In Appendix Figure B5 and Appendix Figure B6, we confirm that our findings are robust for the influential citation count which can be retrieved from Semantic Scholar [34]. This consistency across multiple factors underscores the inherent bias of LLMs towards generating references to highly cited papers, irrespective of other characteristics of the references.
5 Large Language Models and Human Citation Networks
@@
Figure 5 displays the properties of the ground truth and generated citation networks. In Figure 5a, we identify the focal paper (in blue), generated references that appear in the introduction (in green) or later in the paper (in yellow), generated references that are linked to ground truth or other generated references (in orange), generated references that are completely isolated (in purple), and ground truth references that are not cited by GPT-4 (in gray). The majority of generated references ($+50\%$) is non-isolated, i.e., linked to the ground truth or generated references but not present in the focal paper itself, followed by a substantial amount of generated references appearing in the introduction and only a small fraction that do not appear in the introduction but still within the focal paper (Figure 5b). The remainder of generated references is completely isolated from the citation network. If GPT-4 did not pick up on human citation patterns, the generated citation network would resemble a random network containing only isolated citations. The heightened citation bias is also most pronounced for references that appear within the introduction or paper, with isolated generated references having the lowest number of citations (Figure 5c). This finding further indicates the tendency for GPT-4 to more easily identify and generate references to highly cited papers. The number of references is similar across all categories, except for the isolated generated references which have substantially less references (Figure 5f).

### v3

@@ Abstract
Citation practices are crucial in shaping the structure of scientific knowledge, yet they are often influenced by contemporary norms and biases. The emergence of Large Language Models (LLMs) introduces a new dynamic to these practices. Interestingly, the characteristics and potential biases of references recommended by LLMs that entirely rely on their parametric knowledge, and not on search or retrieval-augmented generation, remain unexplored. Here, we analyze these characteristics in an experiment using a dataset from AAAI, NeurIPS, ICML, and ICLR, published after GPT-4’s knowledge cut-off date. In our experiment, LLMs are tasked with suggesting scholarly references for the anonymized in-text citations within these papers. Our findings reveal a remarkable similarity between human and LLM citation patterns, but with a more pronounced high citation bias, which persists even after controlling for publication year, title length, number of authors, and venue. The results hold for both GPT-4, and the more capable models GPT-4o and Claude 3.5 where the papers are part of the training data. Additionally, we observe a large consistency between the characteristics of LLM’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. By analyzing citation graphs, we show that the references recommended are embedded in the relevant citation context, suggesting an even deeper conceptual internalization of the citation networks. While LLMs can aid in citation generation, they may also amplify existing biases, such as the Matthew effect, and introduce new ones, potentially skewing scientific knowledge dissemination.
Introduction
@@
Figure 1: Overview of our experiment evaluating the characteristics and biases of LLM generated references, when tasked to suggest references for anonymized in-text citations. We collect $166$ papers from the cs.LG category on arXiv which are published in the main tracks of AAAI, NeurIPS, ICML, and ICLR, and only appeared available online after GPT-4’s knowledge cut-off date. We split the main content, which includes the author information, conference information, abstract, and introduction, from the ground truth references. GPT-4, GPT-4o and Claude 3.5 are prompted to generate suggestions of scholarly references for the anonymized in-text citations in the main content. We verify the existence of the generated references via Semantic Scholar and compare the characteristics, such as title length, publication year, venue, and number of authors, of the existing and non-existent generated references with the ground truth. For the existing generated references, we also compare additional characteristics, such as the number of citations and references, and analyze the properties of their citation networks.
@@
Large Language Models (LLMs) have revolutionized natural language understanding and generation, driving scientific research forward by assisting in all steps of the scientific process, ranging from identifying research gaps to accelerating complex data analysis (Boiko, MacKnight, and Gomes 2023; Merchant et al. 2023; Romera-Paredes et al. 2024; Zheng et al. 2023). One particularly interesting application is the generation of suggestions for appropriate scholarly references (Qureshi et al. 2023; Walters and Wilder 2023). Yet, without the aid of web browsing or retrieval-augmented generation, these models rely entirely on their parametric knowledge encapsulated during their (pre-)training (Brown et al. 2020; Bubeck et al. 2023; Kaddour et al. 2023; Wei et al. 2022a). Our research focuses on this intrinsic citation behavior of GPT-4, exploring how the model recommends references based on its training data, and highlighting the potential biases that arise from this internalized knowledge (Acerbi and Stubbersfield 2023; Manerba et al. 2023).
@@
Biases in citation practices have long been a subject of scrutiny in the scientific community (Fortunato et al. 2018; Smith 2012). Common biases include a preference for recent publications (Bornmann and Daniel 2008), shorter titles (Letchford, Moat, and Preis 2015), and high-profile publication venues (Lawrence 2003). Moreover, a well-documented phenomenon is the “Matthew effect,” where highly cited papers tend to accumulate even more citations (Wang 2014). The number of authors on a paper can also influence its likelihood of being cited, with solo and small author groups often cited less frequently than larger teams (Gazni and Didegah 2011). By examining how these biases manifest in LLM generated references, we aim to uncover the underlying patterns and their potential to amplify existing biases and introduce new ones, potentially skewing scientific knowledge dissemination.
In our experiment, we let GPT-4, GPT-4o and Claude 3.5 suggest scholarly references for anonymized in-text citations within a paper and compare the characteristics and citation networks of the LLM generated references against the ground truth. We provide a comprehensive analysis of $166$ papers which are published in the main tracks of AAAI, NeurIPS, ICML, and ICLR, encompassing 3,066 references in total. All the papers are only first available online on arXiv after GPT-4-0613’s knowledge cut-off date and belong to the cs.LG category. While this experimental setup may not fully reflect real-world usage of LLMs for citation generation, which often involves more interactivity and reliance on external data sources, it provides a controlled laboratory setting to assess the parametric knowledge and inherent biases of LLMs. Furthermore, our focused sample of papers ensures a homogeneous dataset, which allows us to minimize confounding factors that could arise from cross-disciplinary differences in citation practices.
@@
Our setting differs from previous work which either let the LLM generate short papers or literature reviews, or is prompted for the most important papers on a certain topic (Walters and Wilder 2023). We argue that these methods are more susceptible to the LLM’s memorization capabilities (Chen et al. 2024; Kadavath et al. 2022). Moreover, the evaluation of the suggested references mostly focuses on their existence, bibliometric accuracy, or qualitative judgement by domain experts (Qureshi et al. 2023). Finally, another strand of the literature focuses on improving LLMs via search and retrieval-augmented generation (Lewis et al. 2020) or to reduce their hallucination rate via self-consistency (Agrawal et al. 2024) to enhance their capabilities in systematic literature reviews (Susnjak et al. 2024).
In our experiment, we find that GPT-4 exhibits strong preferences for highly cited papers, which persists even after controlling for multiple confounding factors such as publication year, title length, venue, and number of authors. Additionally, we observe a large consistency between GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. The same results hold for the more capable models GPT-4o and Claude 3.5 where the papers are part of the training data. By analyzing citation graphs, we show that the references recommended by GPT-4 are embedded in the relevant citation context, suggesting an even deeper conceptual internalization of the citation networks. While LLMs can aid in citation generation, we find they may also amplify existing biases and introduce new ones, potentially skewing scientific discourse. Our results underscore the need for identifying the model’s biases and for developing balanced methods to interact with LLMs in general (Navigli, Conia, and Ross 2023).
Generating Citations with LLMs
@@
Our data consists of $166$ papers published at AAAI ($25$), NeurIPS ($72$), ICML ($38$), and ICLR ($31$) for a total of $3,066$ references. Our data collection process is depicted in Figure 1 (see Appendix B for more details) and begins by retrieving all the relevant papers from arXiv, focusing on those within the machine learning category (cs.LG) and posted between March 2022 and October 2023 (after GPT-4-0613’s knowledge cut-off date). The papers are verified on Semantic Scholar where we store additional metadata, such as all the reference titles with corresponding Semantic Scholar IDs to construct the citation networks (see Appendix Table LABEL:tab:paper_details for a full list of all the included papers).
@@
We split the main content, which includes the author information, conference information, abstract, and introduction, from the ground truth references. Next, we prompt GPT-4, GPT-4o and Claude 3.5 as follows:
@@
We then post-process the responses to extract the title, venue, publication year, author names, and number of authors for each generated reference (see Appendix B for more details). To assess the robustness of this approach, we repeat this “vanilla” approach three to five times for all $166$ papers.
A well-known issue in text generation by LLMs are hallucinations or confabulations, which refer to generated content that is nonsensical or untruthful in relation to certain sources, i.e., factual mistakes about historical events (Zhang et al. 2023). This is particularly problematic for the generation of scholarly references, as LLMs can fabricate references that do not exist or introduce subtle errors, making it impossible to retrieve the actual references (Walters and Wilder 2023). There are two main approaches to verify the existence of LLM-generated references: one involves asking additional questions to the LLM to verify its self-consistency (Agrawal et al. 2024), and the second approach utilizes external databases to verify a reference’s existence (Fabiano et al. 2024). In our experiment, we opt for the latter and determine via title and author names matching with Semantic Scholar entries whether the generated references exist (see Appendix B for more details). Finally, we also build on our “vanilla” approach, by introducing an “iterative” approach where we continue to prompt GPT-4 after having indicated which generated references do not exist and ask to replace those with existing ones (see Appendix B for more details). The previously existing generated and the newly generated references are then merged.
In Table 1, we report the GPT-4 summary statistics for each of the five vanilla (and iterative: results between brackets) runs. On average, $65\%$ ($86\%$) of the generated references match with an entry in Semantic Scholar, while $13\%$ ($14\%$) and $17\%$ ($20\%$) of them appear in the introduction or paper itself, respectively. We further show that about $7\%$ ($7\%$) of the generated and ground truth references match pairwise. However, this number grows to $13\%$ ($14\%$) if we only consider the uniquely identifiable references (i.e., omitting references included in [4–8] as there is no one-to-one correspondence). In Appendix Table C1, we show that the average overlap between generated sets is $17\%$.
@@
Pairwise Match (PM) 7.0 7.2 6.3 6.9 6.7
for all references (7.1) (7.3) (6.6) (7.0) (7.1)
PM for uniquely 12.5 13.7 12.5 13.7 13.3
@@
Table 1: Summary statistics of GPT-4 generated references. (in % with respect to total number of references).
Figure 2: Properties of the ground truth and GPT-4 generated introduction references for the vanilla strategy. This figure displays the properties of the ground truth ($n=14,554$, in blue) and GPT-4 generated references ($n=14,554$, in green), further subdividing the generated references into existing ($n=9,376$, in orange) and non-existent ($n=5,178$, in red), from the original data sources of five runs for the vanilla strategy with GPT-4. a, The average percentage of existing generated references in total ($64.4\%$) and for each publication venue under the vanilla and iterative strategy, with dots representing the percentage for each of the five runs with GPT-4. b, The distribution of the number of characters in the title shows some differences between the ground truth (median $62$) and generated (median $58$) references with the non-existent being slightly longer and with a larger variance compared to the existing generations. c, The distribution over time reveals that ground truth references are relatively more recent than generated references, with most references post-$2010$. The temporal distribution of the non-existent generated references aligns more with the ground truth than the existing generated references. d, The distribution of the number of authors demonstrates a disparity between the ground truth and generated references, having median values of three and two, respectively. However, GPT-4 more often generates “et al.” which does not allow for an exact computation, especially for the non-existent references. e, The distribution of publication venues shows that for most venues the ground truth has the highest relative representation, followed closely by existing references. The non-existent references deviate more from the ground truth as the proportion of “Others” is substantially larger. f, The distributions of citations for ground truth and existing generated references reveal a substantial citation bias in the generated references with a difference in median citations of $1,326$. g, Finally, the distribution of references shows that ground truth references cite slightly more papers than the existing generated references with a median difference in median references of $6$.
Reflecting Human Citation Patterns
Figure 2 displays the characteristics of the ground truth and GPT-4 generated references, and separately the characteristics of the generated references which match with a Semantic Scholar entry, and those which do not exist according to this database. Overall, we observe a remarkable similarity between human and LLM citation patterns and a large consistency between GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. In Appendix Figure B1, we also show that the newly generated papers from the “iterative” approach show nearly identical distributions.
The distributions of the title lengths show that existing generated reference titles tend to be the shortest, while non-existent generated reference titles are more similar in length to the ground truth, which indicates a learned pattern. Overall, the first effect dominates, so the average is skewed to shorter titles for generated references (Figure 2b). The temporal analysis reveals a similar pattern where non-existent generated references follow a distribution that is more similar to the ground truth than the existent ones (Figure 2c).
@@
Most prominently, we observe a significant citation bias in the existing generated references, which have a median citation count of $1,326$ higher than ground truth references (Figure 2f). In Appendix Figure B6, we compare the characteristics for the corresponding ground truth references of existing and non-existent references, and for the existing references which also appear in the paper itself. We observe that the ground truth papers which correspond to existing references that appear in the paper itself have by far the most citations, followed by the existing references, and the ground truth papers corresponding to non-existent references have the lowest numbers of citations. These findings further indicate the tendency for GPT-4 to more easily generate references to highly cited papers. Finally, the distribution of references indicates that ground truth references cite slightly more papers than existing generated references (Figure 2g).
In Appendix Figures B4 and B5 and Table C2, we find similar results for three GPT-4o and Claude 3.5 runs, but with a higher existence rate which may be due to the models’ capabilities or the papers being part of the training data.
@@
Figure 3: The citation bias in existing GPT-4 generated references is not due to the recency of ground truth references. This figure shows that the existing GPT-4 generated references ($n=9,376$, in orange) consistently exhibit a higher citation count compared to their corresponding ground truth ($n=9,376$, in blue) across subperiods. a, The citation counts across time for the ground truth and existing generated references reveal that the most recent references have a relatively low number of citations. The difference in median citations between the existing generated references and their corresponding ground truth references is 1,257. Since the ground truth references are relatively more recent compared to the existing generated references, we examine whether the observed citation bias is related to the recency of ground truth references. b, The distributions of citations by subperiod reveal that the existing generated references consistently exhibit a higher citation count than their corresponding ground truth counterparts. c, The difference in median citations is most pronounced in the early and late subperiods, i.e., $\leq 1988$, 2010-2016, and 2017-2023.
Heightened Citation Bias
@@
In Appendix Figure B7, we find that the heightened citation bias in generated references remains also after controlling for other possible confounding factors, such as title length, number of authors, and publication venue. In Appendix Figure B8 and Appendix Figure B9, we confirm that our findings are robust for the influential citation count which can be retrieved from Semantic Scholar (Valenzuela-Escarcega, Ha, and Etzioni 2015). This consistency across multiple factors underscores the inherent bias of LLMs towards generating references to highly cited papers, irrespective of other characteristics of the references.
Figure 4: The GPT-4 generated references display similar citation network properties as the ground truth references but with a heightened citation bias. This figure displays how the existing GPT-4 generated references ($n=2945$, first run of vanilla strategy) are embedded in the citation network of the focal papers ($m=166$ in total). a, We depict the connections between the focal paper, the ground truth references, and the existing generated references by showing the underlying citation graphs. An arrow from A to B indicates that A cites B. We identify the focal paper (in blue), generated references that appear in the introduction (in green) or in the paper (in yellow), generated references that are linked to ground truth or other generated references (in orange), generated references that are completely isolated (in purple), and ground truth references that are not cited by GPT-4 (in gray). b, The majority of generated references does not appear in the introduction or paper itself, but is somehow connected to the ground truth references as only a small fraction of generated references is completely isolated. c, The heightened citation bias is most emphasized for generated references that appear in the introduction or the paper, with isolated generated references having the lowest number of citations. d, The normalized average clustering coefficients of the ground truth (green and grey nodes) and the existing generated references (green, yellow, orange, and purple nodes) indicate that GPT-4’s internalization of citation patterns extends to citation network properties. The clustering coefficient for a node A is given by #triangles through A#possible triangles through $\frac{\#\text{triangles through A}}{\#\text{possible triangles through A}}$. The average is computed across the coefficients of all nodes in the respective graph (excl. nodes with coefficient zero) and indicates the tendency of the respective references to appear in clusters. e, The non-isolated generated references are tightly connected to the ground truth references, both on an individual level (Boolean edge density) as well as an aggregate level (edge expansion). The Boolean edge density is the fraction of non-isolated generations (orange nodes) that are connected to at least one ground truth reference (green and grey nodes) per focal paper. The edge expansion between those two sets is defined as the number of edges between the two sets divided by the smallest set size. f, The number of references is similar across all categories, except for the isolated generated references which have substantially less references.
LLMs and Human Citation Networks
@@
Figure 4 displays the properties of the ground truth and GPT-4 generated citation networks. In Figure 4a, we identify the focal paper (in blue), generated references that appear in the introduction (in green) or later in the paper (in yellow), generated references that are linked to ground truth or other generated references (in orange), generated references that are completely isolated (in purple), and ground truth references that are not cited by GPT-4 (in gray). The majority of generated references ($>50\%$) is non-isolated, i.e., linked to the ground truth or generated references but not present in the focal paper itself, followed by a substantial amount of generated references appearing in the introduction and only a small fraction that do not appear in the introduction but still within the focal paper (Figure 4b). The remainder of generated references is completely isolated from the citation network. If GPT-4 did not pick up on human citation patterns, the generated citation network would resemble a random network containing only isolated citations. The heightened citation bias is also most pronounced for references that appear within the introduction or paper, with isolated generated references having the lowest number of citations (Figure 4c). This finding further indicates the tendency for GPT-4 to more easily identify and generate references to highly cited papers. The number of references is similar across all categories, except for the isolated generated references which have substantially less references (Figure 4f).